f(x)=3x+2 and g(x)=2x-3, what is the value of g[f(-2)]

Answer by jim_thompson5910(35256)

f(x) = 3x+2


f(-2) = 3(-2)+2


f(-2) = -6 + 2


f(-2) = -4


-------------------------------------------------------


g(x) = 2x-3


g(f(-2)) = 2(f(-2))-3


g(f(-2)) = 2(-4)-3


g(f(-2)) = -8-3


g(f(-2)) = -11